OpenStudy (anonymous):

Find the equation of the line tangent to the graph of the given function at the point with the indicated x-coordinate. f(x) = ((sqr x)+1)/(sqr x)+3); x=4

OpenStudy (anonymous):

\[f(x)=\frac{ \sqrt{x} +1 }{ \sqrt{x}+3 }; x=4\]